# Towards Holistic Evaluation of Large Audio-Language Models: A Comprehensive Survey

### v3

@@ Abstract
Chih-Kai Yang1, Neo S. Ho, Hung-yi Lee2 National Taiwan University 1chihkaiyang1124@gmail.com, 2hungyilee@ntu.edu.tw
@@ Acknowledgments
We thank the reviewers for their valuable feedback, which have helped us improve the paper. To enhance the comprehensiveness of our survey, we have added further discussion in the appendix.
@@ Appendix B Brief Summary for Benchmarks Discussed in the Main Text
| Auditory Tasks | Common Datasets | |||||
| Audio Tasks | ||||||
| Audio Captioning |
| Audio Classification |
| Vocal Sound Classification | VocalSound (Gong et al., 2022) | |||||
| Speech Tasks | ||||||
| Automatic Speech Recognition |
| Speaker Identification |
| Text-to-Speech |
| Speech Emotion Recognition |
| Language Identification |
| Speech Translation |
| Speech Diarization |
| Keyword Spotting |
| Music Tasks | ||||||
| Music Transcription |
| Instrument Classification | NSynth (Engel et al., 2017) | |||||
| Genre Classification |
| Dataset |
| AudioCaps | 6 | ||
| Clotho | 5 | ||
| ESC-50 | 3 | ||
| AudioSet | 8 | ||
| VocalSound | 2 | ||
| LibriSpeech | 7 | ||
| Common Voice | 8 | ||
| VoxCeleb (1&2) | 3 | ||
| LJSpeech | 3 | ||
| VCTK | 4 | ||
| LibriTTS | 2 | ||
| IEMOCAP | 4 | ||
| CREMA-D | 2 | ||
| VoxLingua107 | 1 | ||
| CoVoST 2 | 2 | ||
| LibriMix | 1 | ||
| Speech Command | 2 | ||
| MusicCaps | 4 | ||
| Song Describer Dataset | 3 | ||
| MAESTRO | 1 | ||
| NSynth | 2 | ||
@@ Appendix H Comparison between End-to-end LALMs and Cascaded Systems
| Type | Benchmark List | ||
Dynamic-SUPERB (Huang et al., 2024a) | ||
| MMAU (Sakshi et al., 2025) | |||
| VoiceBench (Chen et al., 2024c) | |||
| VoxEval (Cui et al., 2025) | |||
| SAKURA (Yang et al., 2025a) | |||
| SD-Eval (Ao et al., 2024) | |||
| Speech-IFEval (Lu et al., 2025) | |||
| Yang et al. (2025b) | |||
SALMon (Maimon et al., 2025) | ||
| Dynamic-SUPERB Phase-2 (Huang et al., 2025a)1 | |||
| AIR-Bench (Yang et al., 2024c)2 | |||
@@ Appendix I Common Metrics for LALMs
| General Auditory Awareness and Processing | ||||||||||
| Benchmark | Input Modalities | Output Modalities | ||||||||
| Text | Audio | Speech | Music | Text | Audio | Speech | Music | |||
| SALMon Maimon et al. (2025) | ✓ | ✓ |
| Wu et al. (2024b) | ✓ | ✓ | ✓ | |||||||
| EmphAssess (Seyssel et al., 2024) | ✓ | ✓ | ||||||||
| Deshmukh et al. (2025b) | ✓ | ✓ | ✓ | ✓ | ||||||
| Dynamic-SUPERB (Huang et al., 2024a) | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
| Dynamic-SUPERB Phase-2 (Huang et al., 2025a) | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
| AIR-Bench (Yang et al., 2024c) | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
| AudioBench (Wang et al., 2025a) | ✓ | ✓ | ✓ | ✓ | ||||||
| MuChoMusic (Weck et al., 2024) | ✓ | ✓ | ✓ | |||||||
| FinAudio (Cao et al., 2025) | ✓ | ✓ | ✓ | |||||||
| SAGI (Bu et al., 2024) | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
| MAE (Chen et al., 2024b) | ✓ | ✓ | ✓ | ✓ | ||||||
| RUListening (Zang et al., 2025) | ✓ | ✓ | ✓ | |||||||
| OpenMU-Bench (Zhao et al., 2024) | ✓ | ✓ | ✓ | |||||||
| Wang et al. (2025b) | ✓ | ✓ | ✓ | |||||||
| Gong et al. (2024a) | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
| Audio-FLAN (Xue et al., 2025) | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ||
| QualiSpeech (Wang et al., 2025c) | ✓ | ✓ | ✓ | |||||||
| EvalSIFT (Pandey et al., 2025) | ✓ | ✓ | ✓ | ✓ | ||||||
| OpenAQA (Gong et al., 2024b) | ✓ | ✓ | ✓ | |||||||
| Clotho-AQA (Lipping et al., 2022) | ✓ | ✓ | ✓ | |||||||
| SpeechCaps (Huang et al., 2025b) | ✓ | ✓ | ✓ | |||||||
| ASR-EC (Wei et al., 2024) | ✓ | ✓ | ✓ | |||||||
| SLU-GLUE (Li et al., 2024) | ✓ | ✓ | ✓ | |||||||
| BEANS-Zero (Robinson et al., 2025) | ✓ | ✓ | ✓ | |||||||
| Knowledge and Reasoning | ||||||||||
| Benchmark | Input Modalities | Output Modalities | ||||||||
| Text | Audio | Speech | Music | Text | Audio | Speech | Music | |||
| ZeroSpeech 2021 (Nguyen et al., 2020) | ✓ |
| CSZS (Huang et al., 2024b) | ✓ |
| sStoryCloze (Hassid et al., 2023) | ✓ |
| tStoryCloze (Hassid et al., 2023) | ✓ |
| BabySLM (Lavechin et al., 2023) | ✓ | ✓ |
| CompA (Ghosh et al., 2024b) | ✓ | ✓ |
| MMAU (Sakshi et al., 2025) | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
| Audiopedia (Penamakuri et al., 2025) | ✓ | ✓ | ✓ | |||||||
| VoiceBench (Chen et al., 2024c) | ✓ | ✓ | ✓ | |||||||
| VoxEval (Cui et al., 2025) | ✓ | ✓ | ||||||||
| SAKURA (Yang et al., 2025a) | ✓ | ✓ | ✓ | ✓ | ||||||
| URO-Bench (Yan et al., 2025) | ✓ | ✓ | ✓ | ✓ | ||||||
| Audio Entailment (Deshmukh et al., 2025a) | ✓ | ✓ | ✓ | |||||||
| ADU-Bench (Gao et al., 2024) | ✓ | ✓ | ✓ | |||||||
| SAGI (Bu et al., 2024) | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
| MuChoMusic (Weck et al., 2024) | ✓ | ✓ | ✓ | |||||||
| RUListening (Zang et al., 2025) | ✓ | ✓ | ✓ | |||||||
| OpenMU-Bench (Zhao et al., 2024) | ✓ | ✓ | ✓ | |||||||
| Gong et al. (2024a) | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
| CompA-R (Ghosh et al., 2024a) | ✓ | ✓ | ✓ | |||||||
| OpenAQA (Gong et al., 2024b) | ✓ | ✓ | ✓ | |||||||
| Clotho-AQA (Lipping et al., 2022) | ✓ | ✓ | ✓ | |||||||
| SLU-GLUE (Li et al., 2024) | ✓ | ✓ | ✓ | |||||||
| SpeechCaps (Huang et al., 2025b) | ✓ | ✓ | ✓ | |||||||
| Wang et al. (2025d) | ✓ | ✓ | ✓ | ✓ | ||||||
| Deshmukh et al. (2025b) | ✓ | ✓ | ✓ | ✓ | ||||||

### v4

@@ Abstract
Chih-Kai Yang1, Neo S. Ho1, Hung-yi Lee1,2 1National Taiwan University, 2NTU Artificial Intelligence Center of Research Excellence (NTU AI-CoRE) chihkaiyang1124@gmail.com https://github.com/ckyang1124/LALM-Evaluation-Survey
@@ Acknowledgments
We thank the reviewers for their valuable feedback, which have helped us improve the paper. To enhance the comprehensiveness of our survey, we have added further discussion in the appendix. This work was supported by the Ministry of Education (MOE) of Taiwan under the project Taiwan Centers of Excellence in Artificial Intelligence, through the NTU Artificial Intelligence Center of Research Excellence (NTU AI-CoRE)
@@ Appendix B Brief Summary for Benchmarks Discussed in the Main Text
| Auditory Tasks | Common Datasets |
| Audio Tasks | |
| Audio Captioning | AudioCaps Kim et al. (2019) Clotho Drossos et al. (2020) |
| Audio Classification | ESC-50 Piczak (2015b) AudioSet Gemmeke et al. (2017) |
| Vocal Sound Classification | VocalSound (Gong et al., 2022) |
| Speech Tasks | |
| Automatic Speech Recognition | LibriSpeech Panayotov et al. (2015b) AISHELL-1 Bu et al. (2017) Common Voice Ardila et al. (2020) |
| Speaker Identification | VoxCeleb2 Chung et al. (2018) CN-Celeb Fan et al. (2020) |
| Text-to-Speech | LJSpeech Ito and Johnson (2017) VCTK Veaux et al. (2017) LibriTTS Zen et al. (2019) |
| Speech Emotion Recognition | IEMOCAP Busso et al. (2008) CREMA-D Cao et al. (2014) |
| Language Identification | VoxLingua107 Valk and Alumäe (2021) FLEURS Conneau et al. (2023) |
| Speech Translation | CoVoST 2 Wang et al. (2021) MuST-C Di Gangi et al. (2019) |
| Speech Diarization | LibriMix Cosentino et al. (2020) |
| Keyword Spotting | Speech Command Warden (2018) |
| Music Tasks | |
| Music Captioning Text-to-Music | MusicCaps Agostinelli et al. (2023) Song Describer Dataset Manco et al. (2023) MidiCaps Melechovsky et al. (2024) |
| Music Transcription | MAESTRO Hawthorne et al. (2019) |
| Instrument Classification | NSynth (Engel et al., 2017) |
| Genre Classification | FMA (Defferrard et al., 2017) GTZAN (Tzanetakis and Cook, 2002) |
| Dataset | # of Benchmarks Using the Dataset |
| AudioCaps | 6 |
| Clotho | 5 |
| ESC-50 | 3 |
| AudioSet | 8 |
| VocalSound | 2 |
| LibriSpeech | 7 |
| Common Voice | 8 |
| VoxCeleb (1&2) | 3 |
| LJSpeech | 3 |
| VCTK | 4 |
| LibriTTS | 2 |
| IEMOCAP | 4 |
| CREMA-D | 2 |
| VoxLingua107 | 1 |
| CoVoST 2 | 2 |
| LibriMix | 1 |
| Speech Command | 2 |
| MusicCaps | 4 |
| Song Describer Dataset | 3 |
| MAESTRO | 1 |
| NSynth | 2 |
@@ Appendix H Comparison between End-to-end LALMs and Cascaded Systems
| Type | Benchmark List |
| Cascaded Systems Outperform | Dynamic-SUPERB (Huang et al., 2024a) |
| MMAU (Sakshi et al., 2025) | |
| VoiceBench (Chen et al., 2024c) | |
| VoxEval (Cui et al., 2025) | |
| SAKURA (Yang et al., 2025a) | |
| SD-Eval (Ao et al., 2024) | |
| Speech-IFEval (Lu et al., 2025) | |
| Yang et al. (2025b) | |
| E2E LALMs Outperform | SALMon (Maimon et al., 2025) |
| Dynamic-SUPERB Phase-2 (Huang et al., 2025a)1 | |
| AIR-Bench (Yang et al., 2024c)2 | |
@@ Appendix I Common Metrics for LALMs
| General Auditory Awareness and Processing | ||||||||
| Benchmark | Input Modalities | Output Modalities | ||||||
| Text | Audio | Speech | Music | Text | Audio | Speech | Music | |
| SALMon Maimon et al. (2025) | ✓ | ✓ | Likelihood-based evaluation. No output modality. | |||||
| Wu et al. (2024b) | ✓ | ✓ | ✓ | |||||
| EmphAssess (Seyssel et al., 2024) | ✓ | ✓ | ||||||
| Deshmukh et al. (2025b) | ✓ | ✓ | ✓ | ✓ | ||||
| Dynamic-SUPERB (Huang et al., 2024a) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| Dynamic-SUPERB Phase-2 (Huang et al., 2025a) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| AIR-Bench (Yang et al., 2024c) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| AudioBench (Wang et al., 2025a) | ✓ | ✓ | ✓ | ✓ | ||||
| MuChoMusic (Weck et al., 2024) | ✓ | ✓ | ✓ | |||||
| FinAudio (Cao et al., 2025) | ✓ | ✓ | ✓ | |||||
| SAGI (Bu et al., 2024) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| MAE (Chen et al., 2024b) | ✓ | ✓ | ✓ | ✓ | ||||
| RUListening (Zang et al., 2025) | ✓ | ✓ | ✓ | |||||
| OpenMU-Bench (Zhao et al., 2024) | ✓ | ✓ | ✓ | |||||
| Wang et al. (2025b) | ✓ | ✓ | ✓ | |||||
| Gong et al. (2024a) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| Audio-FLAN (Xue et al., 2025) | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
| QualiSpeech (Wang et al., 2025c) | ✓ | ✓ | ✓ | |||||
| EvalSIFT (Pandey et al., 2025) | ✓ | ✓ | ✓ | ✓ | ||||
| OpenAQA (Gong et al., 2024b) | ✓ | ✓ | ✓ | |||||
| Clotho-AQA (Lipping et al., 2022) | ✓ | ✓ | ✓ | |||||
| SpeechCaps (Huang et al., 2025b) | ✓ | ✓ | ✓ | |||||
| ASR-EC (Wei et al., 2024) | ✓ | ✓ | ✓ | |||||
| SLU-GLUE (Li et al., 2024) | ✓ | ✓ | ✓ | |||||
| BEANS-Zero (Robinson et al., 2025) | ✓ | ✓ | ✓ | |||||
| Knowledge and Reasoning | ||||||||
| Benchmark | Input Modalities | Output Modalities | ||||||
| Text | Audio | Speech | Music | Text | Audio | Speech | Music | |
| ZeroSpeech 2021 (Nguyen et al., 2020) | ✓ | Likelihood-based evaluation. No output modality. | ||||||
| CSZS (Huang et al., 2024b) | ✓ | Likelihood-based evaluation. No output modality. | ||||||
| sStoryCloze (Hassid et al., 2023) | ✓ | Likelihood-based evaluation. No output modality. | ||||||
| tStoryCloze (Hassid et al., 2023) | ✓ | Likelihood-based evaluation. No output modality. | ||||||
| BabySLM (Lavechin et al., 2023) | ✓ | ✓ | Likelihood-based evaluation. No output modality. | |||||
| CompA (Ghosh et al., 2024b) | ✓ | ✓ | Similarity-based evaluation on text and audio inputs. | |||||
| MMAU (Sakshi et al., 2025) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| Audiopedia (Penamakuri et al., 2025) | ✓ | ✓ | ✓ | |||||
| VoiceBench (Chen et al., 2024c) | ✓ | ✓ | ✓ | |||||
| VoxEval (Cui et al., 2025) | ✓ | ✓ | ||||||
| SAKURA (Yang et al., 2025a) | ✓ | ✓ | ✓ | ✓ | ||||
| URO-Bench (Yan et al., 2025) | ✓ | ✓ | ✓ | ✓ | ||||
| Audio Entailment (Deshmukh et al., 2025a) | ✓ | ✓ | ✓ | |||||
| ADU-Bench (Gao et al., 2024) | ✓ | ✓ | ✓ | |||||
| SAGI (Bu et al., 2024) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| MuChoMusic (Weck et al., 2024) | ✓ | ✓ | ✓ | |||||
| RUListening (Zang et al., 2025) | ✓ | ✓ | ✓ | |||||
| OpenMU-Bench (Zhao et al., 2024) | ✓ | ✓ | ✓ | |||||
| Gong et al. (2024a) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| CompA-R (Ghosh et al., 2024a) | ✓ | ✓ | ✓ | |||||
| OpenAQA (Gong et al., 2024b) | ✓ | ✓ | ✓ | |||||
| Clotho-AQA (Lipping et al., 2022) | ✓ | ✓ | ✓ | |||||
| SLU-GLUE (Li et al., 2024) | ✓ | ✓ | ✓ | |||||
| SpeechCaps (Huang et al., 2025b) | ✓ | ✓ | ✓ | |||||
| Wang et al. (2025d) | ✓ | ✓ | ✓ | ✓ | ||||
| Deshmukh et al. (2025b) | ✓ | ✓ | ✓ | ✓ | ||||